OpenStudy (anonymous):

The table below shows a cummulative frequency distribution of runners' age. AGE GROUP TOTAL 20-29 8 20-39 18 20-49 25 20-59 31 20-69 35 According to the table, how many runners are in their forties? a) 25 b)10 c) 7 d) 6

OpenStudy (amistre64):

25-18 = 15-8 = 7 c